# From Data Deluge to Data Curation: A Filtering-WoRA Paradigm for Efficient Text-based Person Search

### v2

@@ Abstract.
conference: Sydney ’25: ACM The Web Conference; 28 April 2025 - 02 May 2025; Sydney, Australiasubmissionid: 24ccs: Information systems Retrieval efficiencyccs: Information systems Retrieval effectivenessccs: Information systems Multimedia and multimodal retrieval
1. Introduction
Compared to traditional image-based person search  (Zhu et al., 2021; Shao et al., 2022; Wang et al., 2022c, b; Zheng and Zheng, 2024), which seeks to retrieve target individuals from a vast array of footage or images across different locations and times, text-based person search locates interested individuals from a pool of candidates based on pedestrian descriptions (Li et al., 2017a). Given that pedestrian image queries may not be available, text-guided person search emerges as an alternative method. The key lies in mining the fine-grained information from images and texts, blending the complexity of natural language processing with the subtle nuances of visual recognition, and establishing their correspondence. By leveraging the ability to understand complex human descriptions and accurately identify and retrieve images of individuals from a camera system, it can be applied to broad applications in public safety domains such as missing person searches (Syed Musthafa et al., 2024; Ahmad et al., 2023), and rescue operations (Sambolek and Ivasic-Kos, 2021). As the sub-task of vision-language retrieval, text-based person search models usually require extensive data for training, where, however, the number of pedestrian data is limited.
@@ 1. Introduction
  • We introduce a new Filtering-WoRA paradigm for efficient text-based person search, which streamlines learning and improves efficiency through focused data curation and targeted parameter updates. The filtering algorithm targets relevant, high-quality synthesized data by assessing cross-modality relevance, while WoRA (Weighted Low-Rank Adaptation) enables lightweight fine-tuning of a minimal set of model parameters.
@@ 2. Related Work
Vision-Language Pre-training. Current VLP research predominantly bifurcates into coarse-grained and fine-grained methodologies. Coarse-grained approaches employ convolutional networks (Jiang et al., 2020; Huang et al., 2020, 2021; Zhang et al., 2023) or visual Transformers (Kim et al., 2021; Liu et al., 2021; Li et al., 2021; Peng et al., [n. d.]; Kan et al., 2024) to extract and encode holistic image features, thereby constructing vision-language models (VLMs). Techniques such as SOHO (Huang et al., 2021) propose leveraging a Visual Dictionary (VD) for the extraction of comprehensive yet compact image features, facilitating enhanced cross-modal comprehension. ALBEF (Li et al., 2021) introduces a contrastive loss for aligning image and text representations before their fusion through cross-modal attention, fostering more grounded vision-language representation learning. Additionally, it utilizes momentum distillation to augment learning capabilities from noisy network data. Although these holistic image-focused methods are efficient, their performance is generally outpaced by fine-grained approaches. Inspired by advancements in the NLP domain, fine-grained methods (Tan and Bansal, 2019; Lu et al., 2019; Li et al., 2019, 2020a; Chen et al., 2020; Li et al., 2020b; Gan et al., 2020) employ pre-trained object detectors (Ren et al., 2017; Anderson et al., 2018) trained on annotated datasets of common objects, such as COCO (Lin et al., 2014) and Visual Genome (Krishna et al., 2017). This enables models to recognize and classify all potential object regions within images, representing them as a collection of object-centered features. For instance, VinVL (Zhang et al., 2021) enhances visual representations for V+L tasks and develops an improved object detection model to provide object-centered image representations. However, this object-centered feature representation struggles to capture relationships between multiple objects across different regions, limiting its effectiveness in encoding multi-granularity visual concepts. Another limitation is the inability of the object detector to recognize uncommon objects not present in the training data. Recently, novel approaches have emerged to bridge the learning of object-level and image-level alignments. E2E-VLP (Xu et al., 2021) employs DETR (Carion et al., 2020) as the object detection module to enhance detection capabilities. KD-VLP (Liu et al., 2022) relies on external object detectors for object knowledge distillation, facilitating cross-modal alignment learning across different semantic layers. OFA (Wang et al., 2022a) formulates visual-linguistic tasks as a sequence-to-sequence (seq2seq) problem, adhering to instruction-based learning during both pre-training and fine-tuning phases, eliminating the need for extra task-specific layers. Uni-Perceiver (Zhu et al., 2022) constructs a unified perception architecture, using a single Transformer and shared parameters for diverse modes and tasks, employing a non-mixed sampling strategy for stable multi-task learning. X-VLM (Zeng et al., 2022a) and X2-VLM (Zeng et al., 2022b) propose an integrated model with a flexible modular architecture to simultaneously learn multi-granularity alignment and localization, achieving the capability to learn infinite visual concepts related to various text descriptions. In this work, we leverage the proficient vision-language pre-trained model to filter noisy data.
@@
Text-Image Person Search. Based on the challenging task of language-based person search, which is a fine-grained, cross-modal retrieval challenge, a significant number of methodologies have been developed in recent years to tackle this issue. Existing approaches can broadly be classified into two categories: those based on cross-modal attention interaction (Li et al., 2017b; Shao et al., 2022; Shu et al., 2022; Wang et al., 2022c) and those without cross-modal attention interaction (Chen et al., 2022; Ding et al., 2021; Wang et al., 2022b; Zheng et al., 2020b; Yang et al., 2023). Methods leveraging cross-modal attention interactions facilitate correspondences between image regions and textual phrases by pairing inputs and predicting image-text matching scores through attention mechanisms. This enhances interaction between modalities, effectively bridging the modality gap, albeit at the cost of increased computational complexity. For instance, Li et al. (Li et al., 2017b) propose a recurrent neural network with gated neural attention to enhance cross-modal learning. Shao et al. (Shao et al., 2022) introduce a multimodal shared dictionary (MSD) to reconstruct visual and textual features, employing shared, learnable archetypes as queries. To improve person search performance, features for both modalities are extracted in the feature space with uniform granularity, ensuring semantic consistency. Conversely, methods without cross-modal attention interaction, through the construction of diverse model structures and objective functions, align representations of the two modalities within a shared feature space. These lightweight models, not reliant on complex cross-modal interactions, are computationally more efficient and have even achieved better results than their attention-based counterparts. Zheng et al. (Zheng et al., 2020b) build an end-to-end dual-path convolutional network to learn image and text representations to take full advantage of supervision capabilities. However, all the person search methods mentioned above fine-tune the entire network for high accuracy, which inherently slows down the process. In contrast, we propose a method based on cross-modal feature extraction for rapid candidate selection and data ranking scores. This approach aims to mitigate the impact of low-quality text-image pairs while reducing model computational parameters to maintain high efficiency in processing.
@@
Data-Centric Learning. With the surge in popularity of large language models, an increasing demand for vast datasets for model training has emerged (Radford et al., 2021; Zeng et al., 2022a, b; Yang et al., 2023). However, open-source datasets constructed for training models in real-world scenarios, such as the MALS dataset (Yang et al., 2023), may encounter issues like incorrect text descriptions, poor image or text quality, and insufficient feature matching between image-text pairs, all of which can adversely affect model training performance. As the size of datasets expands, it is observed that their quality does not invariably scale in tandem (Zheng et al., 2017). Frequently, a subset of high-caliber data can attain or even exceed the utility of a voluminous but qualitatively inferior dataset. This phenomenon underscores the paramount importance of meticulously curating high-quality datasets, thereby highlighting the necessity for efficiency and precision in dataset construction. For instance, data selection methods (Radenovic et al., 2023; Wu et al., 2024; Huang et al., 2024) aim to identify and train only with the most relevant and informative examples, discarding irrelevant or redundant data. This leads to more efficient learning and improved model performance, especially when dealing with large datasets. Active learning seeks to reduce labeling costs by selecting the most informative instances for annotation. Data cleaning and preprocessing techniques aim to remove noise, errors, and inconsistencies from the data, making it more suitable for learning. Coreset selection (Li et al., 2022, 2023) focuses on identifying a small subset of data points (a core set) sufficient for training a model that performs well across the entire dataset. By selecting a representative subset, coreset selection can reduce the computational costs of training and improve model generalization. Our approach advocates for a coreset method aimed at enhancing model performance, proposing the deploy of the off-the-shelf visual-language models to segment and filter the dataset for text and image pair matching scores, thereby obtaining a coreset dataset for effective training.
@@
Figure 2. An overview of our framework. (a) shows the flow chart of the entire training pipeline. We first obtain the filtered training image-text pairs. Then we augment the text as attribute texts according to keywords. We extract the corresponding features through image encoder, text encoder and cross encoder. There are six loss objectives of both text-image and attribute-image matching tasks. (b) is an in-depth illustration of WoRA methodology, meticulously applied within the context of an image encoder. The model is updated by fine-tuning the decomposition of the pre-trained weights into amplitude and direction components and updating both components using LoRA (Hu et al., 2021) while adding the $\alpha$ and $\beta$ learnable parameters. Since the image encoder consumes most GPU memory and time. In practice, we mainly apply the WoRA on the image encoder.
@@ 3.3. Weighted Low-Rank Adaptation
While the introduction of a pretrain-finetune paradigm to train models for person search tasks has achieved commendable results (Yang et al., 2023), the expansion in model and dataset sizes significantly increases the number of parameters to be trained, demanding substantial computational resources and sacrificing training efficiency. To address this issue, several parameter-efficient fine-tuning methods (Houlsby et al., 2019) have been proposed, aiming to fine-tune pretrained models using the minimum number of parameters. Among these, LoRA (Hu et al., 2021) has gained popularity due to its simplicity and efficacy. LoRA employs a low-rank decomposition for the pretrained weight matrix $W_{0}\in\mathbb{R}^{m\times n}$, $W_{0}+\bigtriangleup W=W_{0}+BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, rank $r\ll\min\left(m,n\right)$. $\bigtriangleup W$ is adjusted by $\frac{8}{r}$. Inspired by the LoRA (Hu et al., 2021), which updates only a small part of the model weight to improve efficiency, DoRA (Liu et al., 2024) decomposes the weight into two parts: direction and amplitude. DoRA improves the adaptability and efficiency of the model, which can be formulated as $W_{DoRA}=m\frac{W_{0}+BA}{\left\|W_{0}+BA\right\|_{2}},$ where $m\in\mathbb{R}^{1\times n}$ is the magnitude vector. $\left\|\cdot\right\|_{2}$ denotes the L2 norm of a matrix across each column.
However, these two methods still limit the parameter freedom. Therefore, we introduce the Weighted Low-Rank Adaptation (WoRA) model to address the capacity gap still present between LoRA (Hu et al., 2021) and fine-tuning (see Figure 2 right). Drawing from the DoRA (Liu et al., 2024) approach, which reparameterizes model weights into magnitude and direction components for fine-tuning, WoRA introduces new learnable parameters, $\alpha$ and $\beta$, to facilitate parameter learning. Given that pretrained weights already possess a vast repository of knowledge suitable for various downstream tasks, we configure these learnable parameters to ensure the model acquires sufficient capability in both magnitude and direction. This allows for the model to adapt to downstream tasks by updating parameters that exhibit significant magnitude or directional changes. The formal representation of our WoRA model is:
@@
In our model, parameters denoted with an overline represent trainable parameters. The proposed Weighted Low-Rank Adaptation (WoRA) demonstrates learning capabilities comparable to full fine-tuning. During inference, WoRA integrates with pretrained weights, introducing no additional latency while enhancing both learnability and computational efficiency. Moreover, by incorporating new learnable parameters, WoRA achieves superior performance compared to LoRA (Hu et al., 2021) and DoRA (Liu et al., 2024).
@@ 4.1. Experimental Setup
LBUL (Wang et al., 2022c) - 64.04 82.66 87.22 -
CAIBC (Wang et al., 2022b) - 64.43 82.87 88.37 -
@@ 4.3. Ablation Study and Further Discussion
Moreover, we assess the impact of Weighted Low-Rank Adaptation (WoRA) on the complete model. Initially, comparing the impacts of using LoRA (Hu et al., 2021) and DoRA (Liu et al., 2024) models of Low-Rank Adaptation on our fully trained model revealed improvements in model speed but not in performance. Subsequently, applying WoRA to models trained on the top50 pre-training selection and the ft90% finetuned dataset, we achieve a Recall@1 of 76.38%, surpassing the baseline by 0.96%, and an mAP of 67.22%, exceeding the baseline by 0.61%, thereby establishing state-of-the-art (SOTA) mAP performance. The computation time for the complete model is reduced from 23h to 18h, marking a 19.82% acceleration. In Figure  7, we compare the effects of different rank values on the performance of WoRA. The experiment shows that the model has the best performance on Recall@1,5,10 when rank is equal to 8. We can observe that model performance may not be as sensitive in different ranks. Performance can be improved basically with the WoRA model. The results we present in this paper are obtained using rank=8, which has achieved the best performance on Recall@1,5,10. As shown in Table  5, we could observe two points (1) WoRA is better than both LoRA and DoRA, whether the performance of Recall@1 or mAP. (2) Furthermore, if we both leverage the top50 and ft90% filtering strategy, WoRA is also the best performing, which surpasses the baseline by 0.71% Recall@1 and 0.95% mAP under the same conditions. Meanwhile, in order to compare the effectiveness of WoRA. We control all the learning rates unchanged and compare the performance of LoRA, DoRA, and WoRA (ous) in Figure  8. we use the same ranks for LoRA, DoRA, and WoRA for comparison. We can clearly see from the figure that under the condition of a fixed learning rate, the WoRA method proposed by us is consistently better than the LoRA and DoRA model mAP under the condition of the same rank. For example, when the same rank is 64, our method is +0.57% mAP higher than LoRA. +2.38% mAP higher than DoRA. Therefore, it can be clearly seen that our WoRA method is superior to DoRA and LoRA under the condition of a fixed learning rate.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714788isbn: 979-8-4007-1274-6/25/04ccs: Information systems Retrieval efficiencyccs: Information systems Retrieval effectivenessccs: Information systems Multimedia and multimodal retrieval
@@
1. Introduction
Compared to traditional image-based person search  (Zhu et al., 2021; Shao et al., 2022; Wang et al., 2022d, c; Zheng and Zheng, 2024; Wang et al., 2023, 2022b), which seeks to retrieve target individuals from a vast array of footage or images across different locations and times, text-based person search locates interested individuals from a pool of candidates based on pedestrian descriptions (Li et al., 2017a). Given that pedestrian image queries may not be available, text-guided person search emerges as an alternative method. The key lies in mining the fine-grained information from images and texts, blending the complexity of natural language processing with the subtle nuances of visual recognition, and establishing their correspondence. By leveraging the ability to understand complex human descriptions and accurately identify and retrieve images of individuals from a camera system, it can be applied to broad applications in public safety domains such as missing person searches (Syed Musthafa et al., 2024; Ahmad et al., 2023), and rescue operations (Sambolek and Ivasic-Kos, 2021). As the sub-task of vision-language retrieval, text-based person search models usually require extensive data for training, where, however, the number of pedestrian data is limited.
@@ 1. Introduction
  • We propose Filtering-WoRA, a paradigm for efficient text-based person search that streamlines learning through focused data curation and targeted parameter updates. Our filtering algorithm assesses cross-modality relevance to select high-quality synthesized data, while Weighted Low-Rank Adaptation (WoRA) enables lightweight fine-tuning of a minimal set of model parameters.
@@ 2. Related Work
Vision-Language Pre-training. Current VLP research predominantly bifurcates into coarse-grained and fine-grained methodologies. Coarse-grained approaches employ convolutional networks (Jiang et al., 2020; Huang et al., 2020, 2021; Zhang et al., 2023) or visual Transformers (Kim et al., 2021; Liu et al., 2021; Li et al., 2021; Peng et al., 2022; Kan et al., 2024) to extract and encode holistic image features, thereby constructing vision-language models (VLMs). Techniques such as SOHO (Huang et al., 2021) propose leveraging a Visual Dictionary (VD) for the extraction of comprehensive yet compact image features, facilitating enhanced cross-modal comprehension. ALBEF (Li et al., 2021) introduces a contrastive loss for aligning image and text representations before their fusion through cross-modal attention, fostering more grounded vision-language representation learning. Additionally, it utilizes momentum distillation to augment learning capabilities from noisy network data. Although these holistic image-focused methods are efficient, their performance is generally outpaced by fine-grained approaches. Inspired by advancements in the NLP domain, fine-grained methods (Tan and Bansal, 2019; Lu et al., 2019; Li et al., 2019, 2020a; Chen et al., 2020; Li et al., 2020b; Gan et al., 2020) employ pre-trained object detectors (Ren et al., 2017; Anderson et al., 2018) trained on annotated datasets of common objects, such as COCO (Lin et al., 2014) and Visual Genome (Krishna et al., 2017). This enables models to recognize and classify all potential object regions within images, representing them as a collection of object-centered features. For instance, VinVL (Zhang et al., 2021) enhances visual representations for V+L tasks and develops an improved object detection model to provide object-centered image representations. However, this object-centered feature representation struggles to capture relationships between multiple objects across different regions, limiting its effectiveness in encoding multi-granularity visual concepts. Another limitation is the inability of the object detector to recognize uncommon objects not present in the training data. Recently, novel approaches have emerged to bridge the learning of object-level and image-level alignments. E2E-VLP (Xu et al., 2021) employs DETR (Carion et al., 2020) as the object detection module to enhance detection capabilities. KD-VLP (Liu et al., 2022) relies on external object detectors for object knowledge distillation, facilitating cross-modal alignment learning across different semantic layers. OFA (Wang et al., 2022a) formulates visual-linguistic tasks as a sequence-to-sequence (seq2seq) problem, adhering to instruction-based learning during both pre-training and fine-tuning phases, eliminating the need for extra task-specific layers. Uni-Perceiver (Zhu et al., 2022) constructs a unified perception architecture, using a single Transformer and shared parameters for diverse modes and tasks, employing a non-mixed sampling strategy for stable multi-task learning. X-VLM (Zeng et al., 2021) and X2-VLM (Zeng et al., 2022) propose an integrated model with a flexible modular architecture to simultaneously learn multi-granularity alignment and localization, achieving the capability to learn infinite visual concepts related to various text descriptions. In this work, we leverage the proficient vision-language pre-trained model to filter noisy data.
@@
Text-Image Person Search. Based on the challenging task of language-based person search, which is a fine-grained, cross-modal retrieval challenge, a significant number of methodologies have been developed in recent years to tackle this issue. Existing approaches can broadly be classified into two categories: those based on cross-modal attention interaction (Li et al., 2017b; Shao et al., 2022; Shu et al., 2022; Wang et al., 2022d) and those without cross-modal attention interaction (Chen et al., 2022; Ding et al., 2021; Wang et al., 2022c; Zheng et al., 2020b; Yang et al., 2024, 2023). Methods leveraging cross-modal attention interactions facilitate correspondences between image regions and textual phrases by pairing inputs and predicting image-text matching scores through attention mechanisms. This enhances interaction between modalities, effectively bridging the modality gap, albeit at the cost of increased computational complexity. For instance, Li et al. (Li et al., 2017b) propose a recurrent neural network with gated neural attention to enhance cross-modal learning. Shao et al. (Shao et al., 2022) introduce a multimodal shared dictionary (MSD) to reconstruct visual and textual features, employing shared, learnable archetypes as queries. To improve person search performance, features for both modalities are extracted in the feature space with uniform granularity, ensuring semantic consistency. Conversely, methods without cross-modal attention interaction, through the construction of diverse model structures and objective functions, align representations of the two modalities within a shared feature space. These lightweight models, not reliant on complex cross-modal interactions, are computationally more efficient and have even achieved better results than their attention-based counterparts. Zheng et al. (Zheng et al., 2020b) build an end-to-end dual-path convolutional network to learn image and text representations to take full advantage of supervision capabilities. However, all the person search methods mentioned above fine-tune the entire network for high accuracy, which inherently slows down the process. In contrast, we propose a method based on cross-modal feature extraction for rapid candidate selection and data ranking scores. This approach aims to mitigate the impact of low-quality text-image pairs while reducing model computational parameters to maintain high efficiency in processing.
@@
Data-Centric Learning. With the surge in popularity of large language models, an increasing demand for vast datasets for model training has emerged (Radford et al., 2021; Zeng et al., 2021, 2022; Yang et al., 2023). However, open-source datasets constructed for training models in real-world scenarios, such as the MALS dataset (Yang et al., 2023), may encounter issues like incorrect text descriptions, poor image or text quality, and insufficient feature matching between image-text pairs, all of which can adversely affect model training performance. As the size of datasets expands, it is observed that their quality does not invariably scale in tandem (Zheng et al., 2017). Frequently, a subset of high-caliber data can attain or even exceed the utility of a voluminous but qualitatively inferior dataset. This phenomenon underscores the paramount importance of meticulously curating high-quality datasets, thereby highlighting the necessity for efficiency and precision in dataset construction. For instance, data selection methods (Radenovic et al., 2023; Wu et al., 2024; Huang et al., 2024) aim to identify and train only with the most relevant and informative examples, discarding irrelevant or redundant data. This leads to more efficient learning and improved model performance, especially when dealing with large datasets. Active learning seeks to reduce labeling costs by selecting the most informative instances for annotation. Data cleaning and preprocessing techniques aim to remove noise, errors, and inconsistencies from the data, making it more suitable for learning. Coreset selection (Li et al., 2022, 2023) focuses on identifying a small subset of data points (a core set) sufficient for training a model that performs well across the entire dataset. By selecting a representative subset, coreset selection can reduce the computational costs of training and improve model generalization. Our approach advocates for a coreset method aimed at enhancing model performance, proposing the deploy of the off-the-shelf visual-language models to segment and filter the dataset for text and image pair matching scores, thereby obtaining a coreset dataset for effective training.
@@
Figure 2. An overview of our framework. (a) shows the flow chart of the entire training pipeline. We first obtain the filtered training image-text pairs. Then we augment the text as attribute texts according to keywords. We extract the corresponding features through image encoder, text encoder and cross encoder. There are six loss objectives of both text-image and attribute-image matching tasks. (b) is an in-depth illustration of WoRA methodology, meticulously applied within the context of an image encoder. The model is updated by fine-tuning the decomposition of the pre-trained weights into amplitude and direction components and updating both components using LoRA (Hu et al., 2022) while adding the $\alpha$ and $\beta$ learnable parameters. Since the image encoder consumes most GPU memory and time. In practice, we mainly apply the WoRA on the image encoder.
@@ 3.3. Weighted Low-Rank Adaptation
While the introduction of a pretrain-finetune paradigm to train models for person search tasks has achieved commendable results (Yang et al., 2023), the expansion in model and dataset sizes significantly increases the number of parameters to be trained, demanding substantial computational resources and sacrificing training efficiency. To address this issue, several parameter-efficient fine-tuning methods (Houlsby et al., 2019) have been proposed, aiming to fine-tune pretrained models using the minimum number of parameters. Among these, LoRA (Hu et al., 2022) has gained popularity due to its simplicity and efficacy. LoRA employs a low-rank decomposition for the pretrained weight matrix $W_{0}\in\mathbb{R}^{m\times n}$, $W_{0}+\bigtriangleup W=W_{0}+BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, rank $r\ll\min\left(m,n\right)$. $\bigtriangleup W$ is adjusted by $\frac{8}{r}$. Inspired by the LoRA (Hu et al., 2022), which updates only a small part of the model weight to improve efficiency, DoRA (Liu et al., 2024) decomposes the weight into two parts: direction and amplitude. DoRA improves the adaptability and efficiency of the model, which can be formulated as $W_{DoRA}=m\frac{W_{0}+BA}{\left\|W_{0}+BA\right\|_{2}},$ where $m\in\mathbb{R}^{1\times n}$ is the magnitude vector. $\left\|\cdot\right\|_{2}$ denotes the L2 norm of a matrix across each column.
However, these two methods still limit the parameter freedom. Therefore, we introduce the Weighted Low-Rank Adaptation (WoRA) model to address the capacity gap still present between LoRA (Hu et al., 2022) and fine-tuning (see Figure 2 right). Drawing from the DoRA (Liu et al., 2024) approach, which reparameterizes model weights into magnitude and direction components for fine-tuning, WoRA introduces new learnable parameters, $\alpha$ and $\beta$, to facilitate parameter learning. Given that pretrained weights already possess a vast repository of knowledge suitable for various downstream tasks, we configure these learnable parameters to ensure the model acquires sufficient capability in both magnitude and direction. This allows for the model to adapt to downstream tasks by updating parameters that exhibit significant magnitude or directional changes. The formal representation of our WoRA model is:
@@
In our model, parameters denoted with an overline represent trainable parameters. The proposed Weighted Low-Rank Adaptation (WoRA) demonstrates learning capabilities comparable to full fine-tuning. During inference, WoRA integrates with pretrained weights, introducing no additional latency while enhancing both learnability and computational efficiency. Moreover, by incorporating new learnable parameters, WoRA achieves superior performance compared to LoRA (Hu et al., 2022) and DoRA (Liu et al., 2024).
@@ 4.1. Experimental Setup
LBUL (Wang et al., 2022d) - 64.04 82.66 87.22 -
CAIBC (Wang et al., 2022c) - 64.43 82.87 88.37 -
@@ 4.3. Ablation Study and Further Discussion
Moreover, we assess the impact of Weighted Low-Rank Adaptation (WoRA) on the complete model. Initially, comparing the impacts of using LoRA (Hu et al., 2022) and DoRA (Liu et al., 2024) models of Low-Rank Adaptation on our fully trained model revealed improvements in model speed but not in performance. Subsequently, applying WoRA to models trained on the top50 pre-training selection and the ft90% finetuned dataset, we achieve a Recall@1 of 76.38%, surpassing the baseline by 0.96%, and an mAP of 67.22%, exceeding the baseline by 0.61%, thereby establishing state-of-the-art (SOTA) mAP performance. The computation time for the complete model is reduced from 23h to 18h, marking a 19.82% acceleration. In Figure  7, we compare the effects of different rank values on the performance of WoRA. The experiment shows that the model has the best performance on Recall@1,5,10 when rank is equal to 8. We can observe that model performance may not be as sensitive in different ranks. Performance can be improved basically with the WoRA model. The results we present in this paper are obtained using rank=8, which has achieved the best performance on Recall@1,5,10. As shown in Table  5, we could observe two points (1) WoRA is better than both LoRA and DoRA, whether the performance of Recall@1 or mAP. (2) Furthermore, if we both leverage the top50 and ft90% filtering strategy, WoRA is also the best performing, which surpasses the baseline by 0.71% Recall@1 and 0.95% mAP under the same conditions. Meanwhile, in order to compare the effectiveness of WoRA. We control all the learning rates unchanged and compare the performance of LoRA, DoRA, and WoRA (ous) in Figure  8. we use the same ranks for LoRA, DoRA, and WoRA for comparison. We can clearly see from the figure that under the condition of a fixed learning rate, the WoRA method proposed by us is consistently better than the LoRA and DoRA model mAP under the condition of the same rank. For example, when the same rank is 64, our method is +0.57% mAP higher than LoRA. +2.38% mAP higher than DoRA. Therefore, it can be clearly seen that our WoRA method is superior to DoRA and LoRA under the condition of a fixed learning rate.